# Electromagnetic Property Sensing in ISAC with Multiple Base Stations: Algorithm, Pilot Design, and Performance Analysis

### v1

@@ III EM Property Sensing
In this section, we propose an EM property sensing method by fusing received pilot signals of all $K$ subcarriers from the same BS. Denote the central angular frequency as $\omega_{c}=2\pi f_{c}$ and denote the corresponding contrast vector as $\boldsymbol{\chi}_{c}=\boldsymbol{\varepsilon}_{r}-\mathbf{1}_{M}+\frac{j\boldsymbol{\sigma}}{\omega_{c}\boldsymbol{\varepsilon}_{0}}$. Denote $\mathbf{z}_{k,l}=\left[\Re\left(\mathbf{y}_{k,l}\right)^{\top},\Im\left(\mathbf{y}_{k,l}\right)^{\top}\right]^{\top}$ and the EM property vector as $\mathbf{s}=\left[(\boldsymbol{\varepsilon}_{r}-\mathbf{1}_{M})^{\top},(\frac{1}{\omega_{c}\boldsymbol{\varepsilon}_{0}}\boldsymbol{\sigma})^{\top}\right]^{\top}$. We then define the real-valued weighted sensing matrix as:
@@
Given that the region $D$ is mainly filled with air and the target occupies a relatively small part within $D$, most of the elements of $\mathbf{s}$ are 00 corresponding to the EM property of the air. Additionally, $(\boldsymbol{\varepsilon}_{r}-\mathbf{1}_{M})$ and $\frac{1}{\omega_{c}\boldsymbol{\varepsilon}_{0}}\boldsymbol{\sigma}$ share the same support set and the same sparse structure. Utilizing the compressive sensing techniques, $\mathbf{s}$ can be reconstructed by the $l$-th BS individually as the solution to the following regularized mixed $\ell(1,2)$-norm minimization problem:
@@ IV Feature-Level Multi-BS Data Fusion
In this section, we design the multi-BS data fusion algorithm to improve the accuracy of RPCD reconstructions compared to the single-BS sensing results. In order to reduce the transmission overhead between BSs and the central processing unit, we perform the feature-level fusion based on the EM property of the collected data. In this approach, only the most relevant features of RPCD are extracted and transmitted, rather than the raw data, i.e., the received signals. Thus, the feature-level fusion reduces the amount of information exchanged while preserving the quality and utility of the data for further processing and analysis.
@@ IV-A Multi-Agent Consensus Equilibrium
In the multi-BS network, each BS collects information from an individual perspective and can be considered as an agent for the shared EM property sensing task. Therefore, we utilize the multi-agent consensus equilibrium (MACE) framework to harmonize the sensing results of multiple BSs [40], with the goal of enhancing the quality of RPCD reconstruction. In order to solve problem (17) from signals received at all BSs, we introduce the general MAP cost function $h(\mathbf{s})$ as the sum of $h_{l}(\mathbf{s})$, aiming to minimize these individual cost functions collaboratively, i.e.,
@@
The core concept of MACE is that each BS applies an abstract force, represented as an offset, to a consensus solution. MACE is achieved when the collective forces exerted by all the BSs balance each other out, which is called equilibrium in physics [41]. Specifically, the consensus solution of all BSs, defined as $\mathbf{s}^{*}\in\mathbb{R}^{M\times 1}$, is found by solving the following set of MACE equations:
@@
where $\mathbf{u}_{l}^{*}\in\mathbb{R}^{M\times 1}$ represents the abstract force applied by the $l$-th BS that should satisfy (21) in order to achieve equilibrium among all BSs.
@@
Note that (23) always holds irrespective of the choice of $\sigma$, which only influences the time to reach consensus. Considering the convexity of $h(\cdot)$, it has been proven that the stationary point $\mathbf{s}^{*}$ is the global minimum of the general MAP cost function $h(\cdot)$. Therefore, the solution $\mathbf{s}^{*}$ to the MACE equations (20) and (21) is exactly the solution to the MAP reconstruction problem of equation (18).
@@
The reason for the above transformation is that the first term in (25a) can be regarded as the fidelity term, while the remaining three terms in (25a) can be regarded as the regularization terms given by prior knowledge. Since both the fidelity term and the regularization terms are convex functions, ADMM guarantees the finding of the global minimum of the objective function. To equivalently convert (25) into an optimization problem without the constraint, the augmented Lagrangian function $L_{l}(\mathbf{v},\mathbf{a},\mathbf{b})$ associated with the optimization problem (25) is defined as
@@
where $\mathbf{b}\in\mathbb{R}^{2M\times 1}$ is the Lagrangian multiplier and $\eta_{l}$ is the penalty factor. Given the initial vector set $\left\{\mathbf{v}^{(0)},\mathbf{a}^{(0)},\mathbf{b}^{(0)}\right\}$ and until a certain stopping criterion is satisfied, ADMM generates the sequence $(\mathbf{v}^{(n)},\mathbf{a}^{(n)},\mathbf{b}^{(n)})$ recursively as follows:
@@
$+\left.\!\lambda_{l}\|\mathbf{a}\|_{1,2}\!+\!\frac{\|\mathbf{a}-\mathbf{s}\|_{2}^{2}}{2\sigma^{2}\zeta_{l}}\!+\!\frac{\|\mathbf{a}-\mathbf{b}^{(n)}-\mathbf{v}^{(n+1)}\|_{2}^{2}}{2\eta_{l}}\!\right\}\!\!,$ (28)
@@
The computational complexity of the operator $F_{l}(\cdot)$ is dominated by the update of $\mathbf{v}^{(n)}$ in ADMM. Suppose the maximum iteration number is $N_{ADMM}$. Then the complexity of calculating $F_{l}(\cdot)$ is $O(N_{ADMM}(UIN_{r}KM^{2}+M^{3}))$ in the worst case.
@@ IV-B Finding MACE by Solving Fixed-Point Problem
where $\mathbf{S}^{*}$ is the solution of (36). Equation (36) can be interpreted as finding a consensus equilibrium among all BSs. The final MACE reconstruction is then given by $\mathbf{s}^{*}=\bar{x}\left(\mathbf{S}^{*}\right)$, and the columns of $\mathbf{F}\left(\mathbf{S}^{*}\right)-\mathbf{S}^{*}$ are $\mathbf{u}_{l}^{*}$ in (21) $\,,l=1,\ldots,L,$ whose summation is $\mathbf{0}_{2M}$.
@@
The non-expansive property of $\mathbf{T}$ is verified as follows. Since each $F_{l}(\cdot)$ is the proximal mapping of a convex function $h_{l}(\cdot)$, $F_{l}(\cdot)$ belongs to the class of resolvents. Therefore, both the concatenated proximal mapping resolvent $\mathbf{F}$ and the reflection resolvent $2\mathbf{F}-\mathbf{I}$ are non-expansive [42]. Besides, it can be easily proven that the operator $2\mathbf{G}-\mathbf{I}$ is equivalent to right-multiplying $\mathbf{Q}^{(k)}$ by the matrix $(\frac{2}{L}\mathbf{1}_{L}\mathbf{1}_{L}^{\top}-\mathbf{I})$, which is the Householder transformation that belongs to orthogonal transformation and does not change the distance between two vectors. Thus, $2\mathbf{G}-\mathbf{I}$ is also non-expansive. Consequently, $\mathbf{T}$, as the composition operator of $2\mathbf{G}-\mathbf{I}$ and $2\mathbf{F}-\mathbf{I}$, is also non-expansive, and thus the iteration (41) is guaranteed to converge.
@@
During the iterative process of EM property sensing, updating $\tilde{\mathbf{E}}_{l}$ and the function $F_{l}(\cdot)$ is crucial to refine EM property estimates. Initially, $\tilde{\mathbf{E}}_{l}$ is estimated using the Born approximation [37, 38], considering the scattering electric field as relatively minor compared to the incident electric field. With the Born approximation, there is
$\mathbf{D}_{k,l}$ =$=\left\{\bar{\mathbf{H}}_{1,k}^{\top}\left[\mathbf{I}+\sum_{i=1}^{\infty}\left((\mathbf{G}_{k}\textrm{diag}\left({\boldsymbol{\chi}_{k}})^{-\top}\right)^{i}\right)\right]\right\}\circ\mathbf{H}_{2,k,l}$
$\approx\bar{\mathbf{H}}_{1,k}^{\top}\circ\mathbf{H}_{2,k,l}.$ (42)
Approximation (42) allows for the calculation of the initial sensing matrix $\tilde{\mathbf{E}}_{l}^{(0)}$, which are then used to reconstruct the initial guess $\mathbf{S}^{(0)}$. In each iteration, $\tilde{\mathbf{E}}_{l}^{(k)}$ is updated to reflect the latest estimates of EM property, and $F_{l}^{(k)}(\cdot)$ is adjusted accordingly to compensate for the errors introduced by approximation (42). This iterative adjustment leverages recursive linear approximations, which makes the RPCD reconstruction converge to accurate EM property of the target.
The proposed iteration procedure of multi-BS data fusion is summarized in Algorithm 1. Note that only $\mathbf{S}^{(k)}$ is transmitted between BSs and the central processing unit while $\tilde{\mathbf{E}}_{l}^{(k)}$ and $F_{l}^{(k)}(\cdot)$ can be updated by each BS locally without transmission to the central processing unit. The BSs only need to transmit a $2M$-dimensional vector $\mathbf{s}_{l}^{(k)}$ to the central processing unit, saving the overhead of transmitting the general sensing matrix $\tilde{\mathbf{E}}_{l}$, whose dimension is $2UIN_{r}K\times 2M$. Thus, the transmission overhead is significantly decreased in Algorithm 1 compared to signal-level data fusion.
@@ V-A Joint Design of Multi-Subcarrier Pilots
To enhance the diversity of incident EM wave patterns within region $D$, we can minimize the column mutual coherence of the equivalent channels from the $u$-th UE to the target, i.e., $\tilde{\mathbf{H}}_{1,k,u}\overset{\Delta}{=}\mathbf{H}_{1,k,u}\tilde{\mathbf{W}}_{k,u}$. The column mutual coherence, which measures the correlation or linear dependence between signals or patterns, is a critical metric in this context [43, 44]. Reducing column mutual coherence enhances the clarity and distinctiveness of each pilot pattern and is crucial for effective EM property sensing. Since UEs typically do not have access to other UEs’ CSI, each UE should independently design its pilot patterns.
Define $\mathbf{A}_{k,u}=\left[\mathbf{H}_{1,1,u}\tilde{\mathbf{W}}_{1,u},\cdots,\mathbf{H}_{1,k-1,u}\tilde{\mathbf{W}}_{k-1,u}\right]$, and define $\mathbf{A}_{1,u}=\mathbf{0}_{M}$ as a special case with $k=1$. For each UE, we design the pilots for each subcarrier sequentially from $k=1$ to $K$ based on the following optimization:
@@
In the objective function (43a), the first term aims at the reduction of the mutual coherence among different symbols transmitted within the same subcarrier. This is achieved by making the Gram matrix of the equivalent channel as orthogonal as possible, for each UE and at each subcarrier. Therefore, the ideal outcome is that the Gram matrix resembles $\xi_{k,u}\mathbf{I}$. The second term in (43a) aims at the reduction of the mutual coherence between symbols transmitted across different subcarriers. By minimizing this norm, the optimization ensures that the pilot patterns in region $D$ across different subcarriers also remain as orthogonal as possible to each other, which is crucial to improve the overall system performance in multi-subcarrier ISAC environments. Therefore, a fully digital beamforming architecture is necessitated to adjust $\tilde{\mathbf{W}}_{k,u}$ at different carriers.

### v2

@@ III EM Property Sensing
In this section, we propose an EM property sensing method by fusing received pilot signals of all $K$ subcarriers from the same BS. Denote the central angular frequency as $\omega_{c}=2\pi f_{c}$ and denote the corresponding contrast vector as $\boldsymbol{\chi}_{c}$. Denote $\mathbf{z}_{k,l}=\left[\Re\left(\mathbf{y}_{k,l}\right)^{\top},\Im\left(\mathbf{y}_{k,l}\right)^{\top}\right]^{\top}$ and the EM property vector as $\mathbf{s}=\left[(\boldsymbol{\varepsilon}_{r}-\mathbf{1}_{M})^{\top},(\frac{1}{\omega_{c}\boldsymbol{\varepsilon}_{0}}\boldsymbol{\sigma})^{\top}\right]^{\top}$. We then define the real-valued weighted sensing matrix as:
@@
Real-world objects with irregular shapes often exhibit sparsity when placed within artificially defined and regularly shaped regions. [^1] Since the BSs are unaware of the target’s accurate size, the BSs need to choose a large enough $D$ to guarantee that the target with the possibly largest size can be contained by $D$. Besides, the region $D$ is in squared shape while the target’s shape is often irregular. Moreover, if there is no high requirement for the detailed shape of the reconstructed target, the range of $D$ can be expanded to include the surrounding air, thereby maintaining the sparse structure on $\mathbf{s}$ at the cost of reducing the resolution of the target’s shape. Thus, some part of the region $D$ is filled with air under most circumstances, which makes the sparse structure on $\mathbf{s}$ holds true. Given that the region $D$ is mainly filled with air and the target occupies a relatively small part within $D$, most of the elements of $\mathbf{s}$ are 00 corresponding to the EM property of the air. Additionally, $(\boldsymbol{\varepsilon}_{r}-\mathbf{1}_{M})$ and $\frac{1}{\omega_{c}\boldsymbol{\varepsilon}_{0}}\boldsymbol{\sigma}$ share the same support set and the same sparse structure. When a target occupies $\tilde{K}$ sampling points in the region $D$, both $(\boldsymbol{\varepsilon}_{r}-\mathbf{1})$ and $\frac{1}{\omega_{c}\boldsymbol{\varepsilon}_{0}}\boldsymbol{\sigma}$ will have $\tilde{K}$ non-zero elements in the same positions. Consequently, the vector $\mathbf{s}$, which combines $(\boldsymbol{\varepsilon}_{r}-\mathbf{1}_{M})$ and $\frac{1}{\omega_{c}\boldsymbol{\varepsilon}_{0}}\boldsymbol{\sigma}$, will be jointly $\tilde{K}$-sparse. Utilizing the compressive sensing techniques, $\mathbf{s}$ can be reconstructed by the $l$-th BS individually as the solution to the following regularized mixed $\ell(1,2)$-norm minimization problem:
@@ IV Feature-Level Multi-BS Data Fusion
In this section, we design the multi-BS data fusion algorithm to improve the accuracy of RPCD reconstructions compared to the single-BS sensing results. In order to reduce the transmission overhead between BSs and the CPU, we perform the feature-level fusion based on the EM property of the collected data. In the feature-level fusion, only the sensing results of RPCD are transmitted between the BSs and the CPU, rather than the raw data, i.e., the received signals. Thus, the feature-level fusion reduces the amount of information exchanged while preserving the quality and utility of the data for further processing and analysis.
@@ IV-A Multi-Agent Consensus Equilibrium
In the multi-BS network, each BS collects information from an individual perspective and can be considered as an agent for the shared EM property sensing task. Since only the sensing results of RPCD are transmitted between the BSs and the CPU, the CPU is unaware of the sensing equations (22). The BSs need to solve (22) distributedly and the CPU needs to achieve consensus among the sensing results of different BSs. Therefore, we utilize the multi-agent consensus equilibrium (MACE) framework to harmonize the sensing results of multiple BSs [43], because it efficiently addresses the challenge of fusing the distributed data with low transmission overhead.
In order to solve problem (22) from signals received at all BSs, we introduce the general MAP cost function $h(\mathbf{s})$ as the sum of $h_{l}(\mathbf{s})$, aiming to minimize these individual cost functions collaboratively, i.e.,
@@
The core concept of MACE is that each BS applies an abstract force, represented as an offset, to a consensus solution. MACE is achieved when the collective forces exerted by all the BSs balance each other out, which is called equilibrium in physics [44]. Specifically, the consensus solution of all BSs, defined as $\mathbf{s}^{*}\in\mathbb{R}^{2M\times 1}$, is found by solving the following set of MACE equations:
@@
where $\mathbf{u}_{l}^{*}\in\mathbb{R}^{2M\times 1}$ represents the abstract force applied by the $l$-th BS that should satisfy (26) in order to achieve equilibrium among all BSs.
Theorem 1.
The solution $\mathbf{s}^{*}$ to the MACE equations (25) and (26) is exactly the solution to the MAP reconstruction problem (23).
Proof.
@@
Note that (28) always holds irrespective of the choice of $\sigma$, which only influences the time to reach consensus. Considering the convexity of $h(\cdot)$, we have proven that the stationary point $\mathbf{s}^{*}$ is the global minimum of the general MAP cost function $h(\cdot)$ in (23). ∎
@@
The reason for the above transformation is that the first term in (30a) can be regarded as the fidelity term, while the remaining three terms in (30a) can be regarded as the regularization terms given by prior knowledge. Since both the fidelity term and the regularization terms are convex functions, ADMM guarantees the convergence to the global minimum of the objective function (30a) [45]. To equivalently convert (30) into an optimization problem without the constraint, the augmented Lagrangian function $L_{l}(\mathbf{v},\mathbf{a},\mathbf{b})$ associated with the optimization problem (30) is defined as
@@
where $\mathbf{b}\in\mathbb{R}^{2M\times 1}$ is the Lagrangian multiplier and $\eta_{l}$ is the penalty factor. Given the initial vector set $\left\{\mathbf{v}^{(0)},\mathbf{a}^{(0)},\mathbf{b}^{(0)}\right\}$ and until a certain stopping criterion is satisfied, ADMM generates the sequence $(\mathbf{v}^{(n)},\mathbf{a}^{(n)},\mathbf{b}^{(n)})$ recursively as follows
@@
$+\left.\!\lambda_{l}\|\mathbf{a}\|_{1,2}\!+\!\frac{\|\mathbf{a}-\mathbf{s}\|_{2}^{2}}{2\sigma^{2}\zeta_{l}}\!+\!\frac{\|\mathbf{a}-\eta_{l}\mathbf{b}^{(n)}-\mathbf{v}^{(n+1)}\|_{2}^{2}}{2\eta_{l}}\!\right\}\!\!,$ (33)
@@
Remark 3.
The computational complexity of calculating $F_{l}(\cdot)$ in (24) is dominated by the update of $\mathbf{v}^{(n)}$ in ADMM. Suppose the maximum iteration number is $N_{ADMM}$. Then the complexity of calculating $F_{l}(\cdot)$ is $O(N_{ADMM}(UIN_{r}KM^{2}+M^{3}))$ in the worst case. However, if we solve (24) through CVX using the default interior point method, then the order of computational complexity is $O(M^{3.5}\log(1/\rho))$ according to [46], where $\rho$ denotes the target tolerance. Since $M$ is typically much larger than $U,I,N_{r}$, and $K$, the proposed ADMM approach enjoys lower computational complexity than that with CVX.
@@ IV-B Finding MACE by Solving Fixed-Point Problem
where $\mathbf{S}^{*}$ is the solution of (41). Equation (41) can be interpreted as finding a consensus equilibrium among all BSs. The final MACE reconstruction is then given by $\mathbf{s}^{*}=\bar{x}\left(\mathbf{S}^{*}\right)$, and the columns of $\mathbf{F}\left(\mathbf{S}^{*}\right)-\mathbf{S}^{*}$ are $\mathbf{u}_{l}^{*}$ in (26), $l=1,\cdots,L,$ whose summation is $\mathbf{0}_{2M}$.
@@
The non-expansive property of $\mathbf{T}$ is verified as follows. Since each $F_{l}(\cdot)$ is the proximal mapping of a convex function $h_{l}(\cdot)$, $F_{l}(\cdot)$ belongs to the class of resolvents. Therefore, both the concatenated proximal mapping resolvent $\mathbf{F}$ and the reflection resolvent $2\mathbf{F}-\mathbf{I}$ are non-expansive [47]. Besides, it can be easily proven that the operator $2\mathbf{G}-\mathbf{I}$ is equivalent to right-multiplying $\mathbf{Q}^{(k)}$ by the matrix $(\frac{2}{L}\mathbf{1}_{L}\mathbf{1}_{L}^{\top}-\mathbf{I})$, which is the Householder transformation that belongs to orthogonal transformation and does not change the distance between two vectors. Thus, $2\mathbf{G}-\mathbf{I}$ is also non-expansive. Consequently, $\mathbf{T}$, as the composition operator of $2\mathbf{G}-\mathbf{I}$ and $2\mathbf{F}-\mathbf{I}$, is also non-expansive, and thus iteration (46) is guaranteed to converge.
@@
Since $\tilde{\mathbf{E}}_{l}$ and $F_{l}(\cdot)$ are related to $\mathbf{s}$, we integrate the Born iterative method (BIM) [30, 42, 33, 31, 32] into MACE. During the iterative process of EM property sensing based on BIM, updating $\tilde{\mathbf{E}}_{l}$ and the function $F_{l}(\cdot)$ is crucial to refine EM property estimates. Initially, $\tilde{\mathbf{E}}_{l}$ is estimated using the Born approximation [40, 41], which assumes that the scattering electric field is relatively minor compared to the incident electric field. With the Born approximation, there is
$\mathbf{D}_{k,l}\!\!$ =$=\!\!\left\{\bar{\mathbf{H}}_{1,k}^{\top}\!\!\left[\mathbf{I}+\sum_{i=1}^{\infty}\left((\mathbf{G}_{k}\textrm{diag}\left({\boldsymbol{\chi}_{k}})^{-\top}\right)^{i}\right)\right]\right\}\circ(\mathbf{P}_{k,l}\mathbf{H}_{2,k,l})$
$\approx\bar{\mathbf{H}}_{1,k}^{\top}\circ(\mathbf{P}_{k,l}\mathbf{H}_{2,k,l}).$ (47)
The Born approximation (47) allows for the calculation of the initial sensing matrix $\tilde{\mathbf{E}}_{l}^{(0)}$, which is then used to reconstruct the initial guess $\mathbf{S}^{(0)}$. In each BIM iteration, $\tilde{\mathbf{E}}_{l}^{(k)}$ is updated to reflect the latest estimates of EM property, and $F_{l}^{(k)}(\cdot)$ is adjusted accordingly to compensate for the errors introduced by approximation (47). This BIM adjustment leverages recursive linear approximations, which makes the RPCD reconstruction converge to accurate EM property of the target [30, 42, 33, 31, 32].
The proposed iteration procedure of multi-BS data fusion based on BIM is summarized in Algorithm 1. Note that only $\mathbf{S}^{(k)}$ is transmitted between BSs and the CPU while $\tilde{\mathbf{E}}_{l}^{(k)}$ and $F_{l}^{(k)}(\cdot)$ can be updated by each BS locally without being transmitted to the CPU. The BSs only need to transmit a $2M$-dimensional vector $\mathbf{s}_{l}^{(k)}$ to the CPU. If it takes $N_{fusion}$ iterations for Algorithm 1 to converge, then the transmission overhead is $2MN_{fusion}$ real-valued numbers. Comparatively, if the BSs adopt signal-level fusion that requires transmitting the general sensing matrix $\tilde{\mathbf{E}}_{l}$, a total of $4UIN_{r}KM$ complex-valued numbers need to be transmitted. Since $2MN_{fusion}<4UIN_{r}KM$ in most cases, the transmission overhead is decreased in Algorithm 1 compared to signal-level data fusion.
@@ V-A Joint Design of Multi-Subcarrier Pilots
Since there is no closed-form expression for the mean squared error (MSE) of RPCD reconstruction, we use mutual coherence as an alternative pilot design criterion. In the context of compressive sensing, mutual coherence, which measures the correlation or linear dependence between signals or patterns, is a crucial metric used to evaluate the accuracy of RPCD reconstruction [48, 49]. By minimizing the column mutual coherence of the equivalent channels from the $u$-th UE to the target, i.e., $\tilde{\mathbf{H}}_{1,k,u}\overset{\Delta}{=}\mathbf{H}_{1,k,u}\tilde{\mathbf{W}}_{k,u}$, we can enhance the diversity of incident EM wave patterns within region $D$. With lower mutual coherence and more diverse EM wave patterns, richer information of the target’s EM property can be extracted, which is crucial for effective EM property sensing. Since the UE typically does not have access to the other UEs’ CSI, each UE should independently design its pilot patterns.
Define $\mathbf{A}_{k,u}=\left[\mathbf{H}_{1,1,u}\tilde{\mathbf{W}}_{1,u},\cdots,\mathbf{H}_{1,k-1,u}\tilde{\mathbf{W}}_{k-1,u}\right]$, and define $\mathbf{A}_{1,u}=\mathbf{0}_{M}$ as a special case for $k=1$. For each UE, we sequentially design the pilots for the $k$-th subcarrier from $1$ to $K$ based on the following optimization:
@@
In the objective function (48a), the first term aims at the reduction of the mutual coherence among different symbols transmitted within the same subcarrier. This is achieved by making the Gram matrix of the equivalent channel as orthogonal as possible, for each UE and on each subcarrier. Therefore, the ideal outcome is that the Gram matrix resembles $\xi_{k,u}\mathbf{I}$. The second term in (48a) aims at the reduction of the mutual coherence between symbols transmitted across different subcarriers. By minimizing this norm, the optimization ensures that the pilot patterns in region $D$ across different subcarriers also remain as orthogonal as possible to each other, which is crucial to improve the overall system performance in multi-subcarrier ISAC environments. Therefore, a fully digital beamforming architecture is necessitated to adjust $\tilde{\mathbf{W}}_{k,u}$ on different carriers.
@@
Theorem 2.
The PGD algorithm applied to problem (48) is guaranteed to converge to a stationary point, which can be either a local minimum or a saddle point.
Proof.
The proof consists of establishing 4 key properties and then applying the convergence theorem for PGD algorithm on problem (48).
1. Continuous differentiability of $p(\tilde{\mathbf{W}}_{k,u},\xi_{k,u})$: The objective function is a multivariate polynomial and is thus continuously differentiable.
2. Lipschitz continuity of the gradient of $p(\tilde{\mathbf{W}}_{k,u},\xi_{k,u})$: The gradient of $p(\tilde{\mathbf{W}}_{k,u},\xi_{k,u})$ is also a multivariate polynomial. Since $\mathbf{W}_{k,u}$ is finite due to the constraints and $\xi_{k,u}$ is also finite according to (52), the gradient is Lipschitz continuous on the constraint set.
3. Closed and convex constraint set:
  1. 1.
    The constraint $\left\|\tilde{\mathbf{W}}_{k,u}\right\|_{F}^{2}\leq P_{k,u}$ defines a closed and convex set (a ball in the Frobenius norm).
  2. 2.
    The constraint $I\xi_{k,u}\geq\bar{P}_{k,u}$ defines a closed halfspace, which is convex.
4. Lower boundedness of (48a) on the constraint set: The objective function (48a) consists of a sum of squared norms and is bounded from below by zero.
Given these 4 properties, we can apply the convergence theorem for PGD on closed and convex constraint set [51]:
Suppose the step size $\alpha_{k,u}$ satisfies $0<\alpha_{k,u}<\frac{1}{L^{\prime}}$, where $L^{\prime}$ is the Lipschitz constant of $\frac{\partial p(\tilde{\mathbf{W}}_{k,u}^{(t)},\xi_{k,u}^{(t)})}{\partial\tilde{\mathbf{W}}_{k,u}^{*}}$. Then:
  1. 1.
    The sequence $\{p(\tilde{\mathbf{W}}_{k,u}^{(t)},\xi_{k,u}^{(t)})\}$ converges when $t\rightarrow\infty$.
  2. 2.
    The sequence $\{\tilde{\mathbf{W}}_{k,u}^{(t)},\xi_{k,u}^{(t)}\}$ converges when $t\rightarrow\infty$.
  3. 3.
    The limit point of $\{\tilde{\mathbf{W}}_{k,u}^{(t)},\xi_{k,u}^{(t)}\}$ is a stationary point of (48) when $t\rightarrow\infty$.
Therefore, the PGD algorithm applied to problem (48) will converge to a stationary point, which can be either a local minimum or a saddle point. ∎
Remark 4.
While Theorem 2 guarantees the convergence to a stationary point, it does not guarantee that this point is a global minimum due to the non-convexity of problem (48). In practice, running the PGD algorithm multiple times with different random initializations is beneficial to increase the likelihood of finding a good local minimum.